# DreamStory: Open-Domain Story Visualization by LLM-Guided Multi-Subject Consistent Diffusion

### v1

@@ 2.2.2 LLM in Image Generation
Advanced visual generation models struggle with low-quality descriptions, which impedes their comprehension of subtle semantics. Many research [72, 73, 74, 75, 76] efforts have aimed to enhance the capabilities of T2I models by refining datasets and modifying user prompts. Specifically, Segalis et al. [72] enhanced generation performance by re-captioning the images using a specialized LLM and retraining a text-to-image model on the updated data. Some approaches propose rewriting user prompts to enhance generated images regarding the aesthetic [75] and NFT market values [76]. Yang et al. [77] utilize language models for planning, recaptioning, and generating images with coherent layouts. Cheng et al. [78] proposes employing LLM to facilitate user editing in an interactive manner.
@@ 3.2 LLM Prompt Generation Model
Inspired by these pioneering works, we designed a prompt generation model based on the CoT strategy for the diffusion model. Our approach simplifies the entire process into a sequence of simple steps: generating prompts for subjects or scenes, annotating whether subjects are present in scenes, and making necessary revisions. Each of these tasks (text understanding or rewriting) is considerably easier due to its widespread presence in the LLM’s training samples compared with that of directly obtaining a suitable prompt for the diffusion model to visualize stories. All the tasks prompts are designed with at least two in-context examples to improve the performance and formatting of the results [83].
@@ 3.3.1 Existing Attention Mechanism
Existing work has verified that self-attention can control the appearance of the generated image [84], while cross-attention controls the layout and can be used to locate the area of the target subject [85].
Based on these discoveries, recent works have verified the appearance information of reference images can be injected into the generation process by substiting [26] or cascading [27] $K$ and $V$ with that of in the reference images. In the multi-subject scenario, however, they failed to keep multi-subject consistency because all subjects in the target image can refer to the information from all reference images regardless of whether their roles are the same.
@@ 3.3.4 Masked Mutual Cross-Attention
Figure 4: Qualitative comparisons of our DreamStory with SOTA approaches on the FSS real story benchmark. Ours, MuDI, and ConsiStory utilize the subject image on the bottom-left as the reference image. In contrast, StoryDiffusion references the subject image on the bottom-right. Different subjects are indicated with different colors. Please visit the project homepage to watch the video.
Figure 5: Qualitative comparisons of our DreamStory with SOTA approaches on the ChatGPT generated story benchmark. Ours, MuDI, and ConsiStory utilize the subject image on the bottom-left as the reference image. In contrast, StoryDiffusion references the subject image on the bottom-right. Different subjects are indicated with different colors. Please visit the project homepage to watch the video.
@@ 4.2.2 Subjective Evaluation
Due to the bias of existing metrics, a user study is conducted to assess subjective results. Given the variability of individual ratings and the broad spectrum of scores across different evaluators, we employed a pairwise comparison in our user study. For each evaluation, two sets of images were randomly displayed, each generated by a different method and accompanied by their respective texts. Participants were asked to judge each metric by selecting one of three options: Image A is superior, Image B is superior, or both are comparable. We engaged 5 independent evaluators for the assessment. Each evaluator conducted 100 reviews per benchmark, culminating in 500 votes in total. The final results were compiled and are presented as percentages.
@@
TABLE III: Quantitative results of ablation study for our DreamStory on the benchmark. Red and blue indicate the best and the second-best performance.
@@
Figure 8: Qualitative comparisons of our DreamStory with StoryGen on their benchmark. All approaches utilize the subject image on the left as the reference image. Different subjects are indicated with different colors. The narrative text of the story is presented below and serves as the input for our DreamStory. Please visit the project homepage to watch the video.
Figure 9: Qualitative comparisons of our DreamStory with TheaterGen on their benchmark. Different subjects are indicated with different colors. The narrative text of the story is presented on the left and serves as the input for our DreamStory. Please visit the project homepage to watch the video.
@@ 4.4.1 Objective Comparison
The overall results are presented in Tab. II. As can be seen from the table, our DreamStory outperforms other methods in terms of all metrics. Notably, the D&C-DS metric of ours is significantly surpassed other methods, exceeding MuDI, ConsiStory, and StoryDiffusion by margins of 0.1034(23.4%), 0.1293(25.1%), and 0.1080(24.7%) respectively in the 2-Subject of DS-500 benchmark. This pattern is mirrored in the 3-Subject benchmark, reinforcing the effectiveness of our method in maintaining multi-subject consistency. The high performance of LLM in annotating subjects is also presented in Tab. IV. Furthermore, our method exhibits a notable advantage on the DS metric, outperforming other training-free SOTA methods (ConsiStory and StoryDiffusion) on the 2-Subject benchmark by at least 0.05 (9.0%). This lead, however, narrows to an approximate average of 0.02 (3.9%) on the 3-Subject benchmark. This is attributed to the limitations of the diffusion model when generating three subjects, leading to a higher likelihood of subject fusion, which is also discussed in MuDI [24].
@@ 4.4.2 Subjective Comparison
Furthermore, we apply our DreamStory to a case from StoryGen [25]. We control the style of the generated results by adding style prompts, e.g., anime style, as shown in Fig. 8. As seen from Fig. 8, our method maintains better consistency while achieving a higher aesthetic appeal. Our training-free approach effectively leverages existing large, high-quality datasets (e.g., LAION-5B [45]), Overcoming the shortcomings of current story visualization datasets, which are smaller and lower quality. Therefore, our method outperforms StoryGen in aesthetics and can be applied to the open-domain where the subjects and styles are considerably diverse. We also conduct the qualitative comparisons with concurrent work TheaterGen [78] on their benchmark, as shown in Fig. 9. Two styles (anime style and real style) are also applied to our DreamStory to show its performance. As we can see from Fig. 9, our DreamStory can generate more aesthetically pleasing images while maintaining better consistency across multiple subjects, such as the clothes of the boy and girl in the figure. These extensive experimental results also prove the advantages of our DreamStory.
@@ 4.5.2 Subjective Comparison
TABLE IV: Quantitative results of LLM accuracy (%) on 400 synthetic benchmark
0-Subjects 1-Subjects 2-Subject 3-Subject
Accuracy 100.00 98.86 95.29 91.28
@@ 4.6 Limitations and Failure Cases
Our method relies on the abilities of both the LLM and diffusion model. Firstly, LLM may have hallucinations when labeling whether the subject is in the scene. As shown in Fig. 10, the scene includes a boy, but the LLM failed to identify the boy. In addition, the LLM may not effectively distinguish between subjects with similar descriptions, such as the pirate and pirate captain in the second row of Fig. 10.
Finally, diffusion models suffer from semantic understanding issues, which may be difficult in multi-subject and multi-attribute generation [24]. As shown in Fig. 11, when the diffusion model failed to generate three subjects for the first time, our method also failed to generate three subjects with consistent appearances. Despite these limitations, our framework still has promising potential as individual models evolve and progress.

### v2

@@ 2.2.2 LLM in Image Generation
Advanced visual generation models struggle with low-quality descriptions, which impedes their comprehension of subtle semantics. Many research [72, 73, 74, 75, 76, 77] efforts have aimed to enhance the capabilities of T2I models by refining datasets and modifying user prompts. Specifically, Segalis et al. [72] enhanced generation performance by re-captioning the images using a specialized LLM and retraining a text-to-image model on the updated data. Some approaches propose rewriting user prompts to enhance generated images regarding the aesthetic [75] and NFT market values [76]. Yang et al. [78] utilizes language models for planning, recaptioning, and generating images with coherent layouts. Cheng et al. [79] proposes employing LLM to facilitate user editing in an interactive manner.
@@ 3.2 LLM Prompt Generation Model
Inspired by these pioneering works, we designed a prompt generation model based on the CoT strategy for the diffusion model. Our approach simplifies the entire process into a sequence of simple steps: generating prompts for subjects or scenes, annotating whether subjects are present in scenes, and making necessary revisions. Each of these tasks (text understanding or rewriting) is considerably easier due to its widespread presence in the LLM’s training samples compared with that of directly obtaining a suitable prompt for the diffusion model to visualize stories. All the task prompts are designed with at least two in-context examples to improve the performance and formatting of the results [84].
@@ 3.3.1 Existing Attention Mechanism
Existing work has verified that self-attention can control the appearance of the generated image [85], while cross-attention controls the layout and can be used to locate the area of the target subject [86]. Based on these discoveries, recent works have verified the appearance information of reference images can be injected into the generation process by substiting [26] or cascading [27] $K$ and $V$ with that of in the reference images. In the multi-subject scenario, however, they failed to keep multi-subject consistency because all subjects in the target image can refer to the information from all reference images regardless of whether their roles are the same.
@@ 3.3.4 Masked Mutual Cross-Attention
Figure 4: Qualitative comparisons of our DreamStory with SOTA approaches on the FSS real story benchmark. Ours, MuDI, and ConsiStory utilize the subject image on the bottom-left as the reference image. In contrast, StoryDiffusion references the subject image on the bottom-right. Different subjects are indicated with different colors. Please refer to the supplementary materials to watch the video.
Figure 5: Qualitative comparisons of our DreamStory with SOTA approaches on the ChatGPT generated story benchmark. Ours, MuDI, and ConsiStory utilize the subject image on the bottom-left as the reference image. In contrast, StoryDiffusion references the subject image on the bottom-right. Different subjects are indicated with different colors. Please refer to the supplementary materials to watch the video.
@@ 4.2.2 Subjective Evaluation
Due to the bias of existing metrics, a user study is conducted to assess subjective results. Given the variability of individual ratings and the broad spectrum of scores across different evaluators, we employed a pairwise comparison in our user study. For each evaluation, two sets of images were randomly displayed, each generated by a different method and accompanied by their respective texts. Participants were asked to judge each metric by selecting one of three options: Image A is superior, Image B is superior, or both are comparable. We engaged 20 independent evaluators for the assessment. Each evaluator conducted 100 reviews per benchmark, culminating in 2000 votes in total. The final results were compiled and are presented as percentages.
@@
TABLE III: Quantitative results of ablation study on the benchmark. Red and blue indicate the best and the second-best performance.
@@
Figure 8: Qualitative comparisons of our DreamStory with StoryGen on their benchmark. All approaches utilize the subject image on the left as the reference image. Different subjects are indicated with different colors. The narrative text of the story is presented below and serves as the input for our DreamStory. Please refer to the supplementary materials to watch the video.
Figure 9: Qualitative comparisons of our DreamStory with TheaterGen on their benchmark. Different subjects are indicated with different colors. The narrative text of the story is presented on the left and serves as the input for our DreamStory. Please refer to the supplementary materials to watch the video.
@@
Figure 12: Qualitative results of our DreamStory with a different random seed. The first two rows display the same set of subjects in real style with different seeds; the last two rows present another set in anime style. Each row contains two subject images on the left and three scene images on the right, which are generated based on the left subjects. Different subjects are indicated with different colors. These results demonstrate the superior consistency and diversity of our DreamStory across various styles.
@@ 4.4.1 Objective Comparison
The overall results are presented in Tab. II. As can be seen from the table, our DreamStory outperforms other methods in terms of all metrics. Notably, the D&C-DS metric of ours is significantly surpassed other methods, exceeding MuDI, ConsiStory, and StoryDiffusion by margins of 0.1034(23.4%), 0.1293(25.1%), and 0.1080(24.7%) respectively in the 2-Subject of DS-500 benchmark. This pattern is mirrored in the 3-Subject benchmark, reinforcing the effectiveness of our method in maintaining multi-subject consistency. Furthermore, our method exhibits a notable advantage on the DS metric, outperforming other training-free SOTA methods (ConsiStory and StoryDiffusion) on the 2-Subject benchmark by at least 0.05 (9.0%). This lead, however, narrows to an approximate average of 0.02 (3.9%) on the 3-Subject benchmark. This is attributed to the limitations of the diffusion model when generating three subjects, leading to a higher likelihood of subject fusion, which is also discussed in MuDI [24].
@@
Figure 14: Qualitative results of our DreamStory with different LLMs (ChatGPT4, LLaMA3-3B, and LLaMA3-1B). The first two rows are the text and images of the subject, and the last two rows are those of the scenes. Different subjects are indicated with different colors. DreamStory generates visually appealing story scenes while maintaining consistency across multiple subjects, demonstrating the robustness of our framework.
@@ 4.4.2 Subjective Comparison
Furthermore, we apply our DreamStory to a case from StoryGen [25]. We control the style of the generated results by adding style prompts, e.g., anime style, as shown in Fig. 8. As seen from Fig. 8, our method maintains better consistency while achieving a higher aesthetic appeal. Our training-free approach effectively leverages existing large, high-quality datasets (e.g., LAION-5B [45]), overcoming the shortcomings of current story visualization datasets, which are smaller and lower quality. Therefore, our method outperforms StoryGen in aesthetics and can be applied to the open-domain where the subjects and styles are considerably diverse. We also conduct the qualitative comparisons with concurrent work TheaterGen [79] on their benchmark, as shown in Fig. 9. Two styles (anime style and real style) are also applied to our DreamStory to show its performance. As we can see from Fig. 9, our DreamStory can generate more aesthetically pleasing images while maintaining better consistency across multiple subjects, such as the clothes of the boy and girl in the figure. We present subject and scene images generated by our DreamStory with different seeds under both real-style and anime-style conditions, as illustrated in Fig. 12. These extensive experimental results also prove the advantages of our DreamStory.
@@ 4.5.1 Objective Comparison
TABLE IV: Quantitative results of LLM accuracy (%) on benchmark
0-Subject 1-Subject 2-Subject 3-Subject
LLaMA3-1B [96] 100.00 98.89 87.44 80.13
LLaMA3-3B [96] 100.00 98.88 91.44 84.67
ChatGPT4 [32] 100.00 98.86 95.29 91.28
@@ 4.5.2 Subjective Comparison
TABLE V: Comparison results of different segmentation models. Similar performance across various segmentation models demonstrates the robustness and efficiency of our DreamStory.
2-Subject 3-Subject
[…]
w/ Grounded-Light-HQSAM [97] 6.71 0.3774 0.6725 0.5406 6.82 0.3792 0.5962 0.2327
w/ Grounded-MobileSAM [98] 6.72 0.3789 0.6707 0.5386 6.83 0.3799 0.5929 0.2284
w/ Grounded-SAM [89] (Ours) 6.72 0.3779 0.6714 0.5444 6.81 0.3791 0.5965 0.2335
Figure 16: Qualitative comparisons of our DreamStory with different segmentation models. Despite variations in masks produced by different SAM models, the final generated scene images remain nearly identical. This demonstrates the robustness of our method, showing its adaptability to different segmentation models.
4.6 Systematic Analysis
In this subsection, we provide a comprehensive analysis to validate the performance and robustness of DreamStory. This includes a series of degradation tests and time-efficiency evaluations, ensuring a thorough assessment of our approach.
4.6.1 Performance with Varying LLMs
To assess the performance of DreamStory, we replaced two smaller LLMs (LLaMA3.2 [96]) and evaluated their accuracy in subject annotation. As shown in Tab.IV, all LLMs achieved over 98% accuracy on the 0-Subject and 1-Subject benchmarks. For the 2-Subject and 3-Subject benchmarks, accuracy decreased slightly with smaller model sizes but remained above 80%. Among the models tested, the largest LLM, ChatGPT4[32], outperformed the others, achieving approximately 95% accuracy for the 2-Subject and 91% for the 3-Subject benchmarks. Additionally, all LLMs successfully generated prompts for subjects and scenes that aligned with the story, likely due to their robust capabilities in text understanding and summarization—well-established tasks within LLM training data. ChatGPT, in particular, excelled at instruction-following, making it the model of choice for our approach. Fig. 14 illustrates final story images generated using different LLMs, further emphasizing the robustness of DreamStory.
4.6.2 Performance with Varying SAMs
We also evaluated DreamStory’s performance with different SAM models, specifically Light-HQSAM [97] and MobileSAM [98]. The results, presented in Tab.V, show that performance degradation with smaller SAMs across various metrics is minimal. Furthermore, Fig.16 displays example masks generated by different SAMs, along with their corresponding final scene images. While smaller SAMs occasionally miss fine details (such as the sleeves of wizard robes in the first two rows), the differences in the final generated images are negligible. These results affirm the robustness and high performance of DreamStory across varying SAM sizes.
4.6.3 Time Efficiency Analysis
The runtime of the DreamStory framework for generating a single story is primarily determined by the time required for the LLM to process requests, which typically takes 3 to 4 minutes, accounting for about 60% of the total time. This time consumption is largely due to the waiting period for network-based API responses, which could be accelerated in future implementations.
In comparison, the image generation phase is faster, averaging 20 to 30 seconds per image. Additionally, the time required for SAM to generate masks is minimal, averaging less than 0.5 seconds, with negligible impact on the overall processing time. Tab. VI summarizes the time required to generate a scene image using different methods. For finetuning-based approaches (e.g., MuDI), approximately 1.5 hours of finetuning is required for 2-subject scenarios, and 2 hours for 3-subject scenarios. In contrast, DreamStory generates each scene image in about 25 seconds, similar to other training-free approaches, highlighting its exceptional time efficiency.
TABLE VI: Average time(s) Consumption on different methods. It includes the time for generating reference subjects and the final scene. Additionally, MUDI requires extra fine-tuning time for each case.
2-Subject 3-Subject
MuDI [24] 5400 7200
ConsiStory [27] 30 38
StoryDiffusion [31] 21 25
DreamStory(Ours) 22 28
@@ 4.7 Limitations and Failure Cases
Our method relies on the abilities of both the LLM and diffusion model. Firstly, LLM may have hallucinations when labeling whether the subject is in the scene. As shown in Fig. 10, the scene includes a boy, but the LLM failed to identify the boy. In addition, the LLM may not effectively distinguish between subjects with similar descriptions, such as the pirate and pirate captain in the second row of Fig. 10. Finally, diffusion models suffer from semantic understanding issues, which may be difficult in multi-subject and multi-attribute generation [24]. As shown in Fig. 11, when the diffusion model failed to generate three subjects for the first time, our method also failed to generate three subjects with consistent appearances. Despite these limitations, our framework still has promising potential as individual models evolve and progress.